OpenStudy (unklerhaukus):

\[(2t^3y-y^3)\text dt-2t^4\text dy=0\]

OpenStudy (unklerhaukus):

\[\frac{\partial M}{\partial y}=2t^3-3y^2\]\[\frac{\partial N}{\partial t}=-8t^3\]\[\frac{\partial M}{\partial y}\neq\frac{\partial N}{\partial t}\] \[R=R(t,y)=t^my^n\]\[\frac{\partial MR(t,y)}{\partial y}=\frac{\partial NR(t,y)}{\partial t}\]\[R\frac{\partial M}{\partial y}+R_yM=R\frac{\partial N}{\partial t}+R_tN\] \[t^my^n(2t^3)+nt^my^{n-1}(2t^3y)=t^my^n(-8t^3)+mt^{m-1}y^n(-2t^4)\]\[2t^{m+3}y^n+2nt^{m+3}y^{n}=-8t^{m+3}y^n-2mt^{m+3}y^n\]\[(10+2n+2m)t^{m+3}y^n=0\]\[(10+2n+2m)=0\]

OpenStudy (unklerhaukus):

how can i find \(n ~~\& ~~m\)

OpenStudy (unklerhaukus):

i need two equations to solve two variables right?

OpenStudy (unklerhaukus):

\[10+2n+2m=0\] \[2n=-2m-10\] \[-2n=2m+10\] \[n=-m-5\] \[-2(-m-5)=2m+10 \] \[2m+10=2m+10\] \[0=0\] like i said Wislar, i need two equations , not just two versions of the same equation

OpenStudy (unklerhaukus):

\[m+3=0\]?

OpenStudy (anonymous):

How m+ 3 = 0 ??

OpenStudy (unklerhaukus):

i dont know, i copied it from the black board

OpenStudy (unklerhaukus):

\[(10+2n+2m)t^{m+3}y^n=0\]

OpenStudy (anonymous):

By the way I know none of it.. In which topic is this ?? Differential or Partial Differential Equations ??

OpenStudy (unklerhaukus):

first order exact equations with an integrating factor a function of both variables